# Ultrafast Solvent Dynamics Drives the Formation of the Hydrated Electron in Photoexcited Water

### v1

@@ 2.3 H Atom Transfer Mechanism
Another frequently observed pathway is the formation of a species in which the dissociating hydrogen atom binds to a water molecule, forming the radical hydronium ($\mathrm{H_{3}O}^{\bullet}$), characterized by $CN_{H}\sim$1 and $CN_{O}\sim$3 (Figure 3C). We confirmed that this species is indeed $\mathrm{H_{3}O}^{\bullet}$ and not the conventional $\mathrm{H_{3}O}^{+}$ ion, since the electron remains associated with it. While the species $\mathrm{H_{3}O}^{\bullet}$ has been predicted by various theoretical models27, 53, 54, 67, our work is the first to demonstrate its formation following excitation by simulating the entire process from photoabsorption to relaxation in liquid water. According to our simulations, the radical $\mathrm{H_{3}O}^{\bullet}$ appears only transiently. It eventually undergoes further dissociation, where the hydrogen atom carrying the electron migrates into a region surrounded by water molecules. In these configurations, an OH bond from a nearby water molecule points toward the hydrogen atom (see Figure 3D) creating an unstable geometry ultimately leading to non-radiative decay to the ground-state. To the best of our knowledge the radical hydronium ($\mathrm{H_{3}O}^{\bullet}$) has not been observed experimentally through photoionization/photoabsorption of water perhaps in part due to the fact that it is formed fleetingly as a rare transition state-like motif.
@@ 2.4 Proton Coupled Electron Transfer Mechanism
Experimental studies have reported initial gyration radius of around 20-40 Å33, 64 immediately after photoabsorption. Although our method captures the same qualitative behavior, direct comparison with experiment is limited by the size of the simulation box (12.42 Å), which restricts the maximum observable extent of delocalization68. The final gyration radius we observe in the excited state is 2.7 Å, slightly larger than the experimentally reported value of 2.4 Å for the hydrated electron in the ground state69. Note that the decrease of the gyration radius from 6 to 2.4 Å incurs a decrease in the energy gap. Once the electron has localized, the system undergoes a non-radiative transition to the ground state, as indicated by the red points in Figure 4A.
As mentioned in the introduction, generating a free electron in bulk water using photon energies within the first absorption band requires some degree of nuclear rearrangement to release the electron from the molecule16. Earlier, we showed that this release occurs simultaneously with the dissociation of the OH bond in the excited water molecule, leading to the formation of a free electron, a hydroxyl radical ($\mathrm{HO}^{\bullet}$), and a proton ($\mathrm{H}^{+}$or $\mathrm{H_{3}O}^{+}$). We now move on to elucidating that the localization of the electron and the formation of the other species requires a collective process involving reorganization of the hydrogen bond network of the water molecules.
In Figure 4B we illustrate the coupling between the gyration radius against the angle formed between the vector connecting the oxygen and the center of the electron and the OH vector of the water molecule. For this analysis we only include the nearby water molecules involved in the first solvation shell of the electron, (see Analysis Protocols section for details). When the electron is delocalized, this angle spans a broad range from 0 to 120 deg, reflecting the absence of a well-defined solvation shell. As the system relaxes, the angle narrows to below 30 deg, consistent with the orientational preferences seen in hydrated electron solvation shells55, 68, 70. These results indicate that the structure of the hydrated electron forms within the excited-state itself, before the non-radiative decay consistent with recent experiments using sub-two-cycle visible-to-shortwave infrared pump–probe spectroscopy71. Analysis of the radial distribution function in the low gyration radius regime on the excited state (see Figure S9 in the Supplementary Information), where the hydrated electron is effectively localized, reveals that the first solvation shell comprises between 4-5 water molecules. This is consistent with previous findings for the hydrated electron in the ground electronic state43, 49, 24.
@@ 2.5 Coupled Translation and Rotational Solvent Dynamics Stabilizes Water-Mediated Photoproducts
In addition to these rotational dynamics, Figure 5B shows that localization also involves collective translational motion. Water molecules in the first solvation shell move on average by approximately 1.2 Å relative to their final positions, a displacement slightly larger than that of the more distant water molecules. Overall, our results show that both rotational and translational rearrangements of specific water molecules are necessary for the localization of the hydrated electron. These collective motions also cause a notable distortion of the hydrogen bond network, consistent with experimental findings from ionization studies in liquid water34.
As alluded to earlier, the creation of the excess electron after photoexcitation also leads to the formation of a hydronium ion ($\mathrm{H_{3}O}^{+}$) and a hydroxyl radical ($\mathrm{HO}^{\bullet}$). To investigate the behavior of these species during the excited-state dynamics, we present in Figure 5C the coupling between the gyration radius of the electron and the distance between the $\mathrm{HO}^{\bullet}$ and the $\mathrm{H_{3}O}^{+}$. Interestingly, our analysis reveals three distinct minima in the $\mathrm{HO}^{\bullet}$($\mathrm{H_{3}O}^{+}$) distance, particularly in the region of low gyration radius where the electron becomes localized. These minima occur at $\sim$2.6, 4.3, and 6.2 Å. Our simulations are in excellent agreement with very recent time-resolved electron diffraction experiments following the ionization of liquid water, where the corresponding minima were found at 2.4, 3.4, and 5.8 Å34. These three minima correspond to situations involving the initial formation of contact ion-radical pair and then the separation mediated by the solvent. Although previous theoretical studies have identified the $\mathrm{HO}^{\bullet}$($\mathrm{H_{3}O}^{+}$) pair by simulating ionized water with the electron removed72, 73, 74, our findings show for the first time, the creation of these correlated structures upon photoexcitation starting from the neutral bulk water.
The creation of ion-radical pairs at different distances has important implications on the optical properties. To probe this, we show in Figure 5D the energy gap as a function of the $\mathrm{HO}^{\bullet}$($\mathrm{H_{3}O}^{+}$) distance. For the contact ion-radical pair at 2.6 Å the average energy gap is $\sim$3 eV on average. As one moves towards the creation of the solvent-separated ion-radical pairs, the energy gap drops to $\sim$2.0 eV. This trend is also consistent with experimental observations, which demonstrated that the formation and spatial separation of the $\mathrm{HO}^{\bullet}$($\mathrm{H_{3}O}^{+}$) pair occur before the internal conversion associated with the $p\to s$ transition of the hydrated electron34.
@@ 2.6 Spontaneous Emission from Hydrated Excess Electrons
An important and still not fully understood property of the hydrated electron in liquid water is its photon emission following excitation. Understanding this emission can provide valuable insight into the nature of the excited state and the relaxation dynamics of the system. Although experimental studies are relatively scarce, Tauber and Mathies have explored this optical behavior through fluorescence and Raman spectroscopies55. Motivated by their findings, we now focus on analyzing the emission spectrum obtained from our simulations and examine how it depends on the gyration radius of the hydrated electron.
Figure 6A shows the experimental emission spectrum along with our calculated $S_{1}\to S_{0}$ energy gaps, evaluated for configurations with both large and small electron gyration radius. As the excess electron becomes more localized, transitioning from a larger to a smaller gyration radius, the energy gap decreases from an average of 2.4 to 1.3 eV (516 to 954 nm). Interestingly, the experimental emission spectrum is in excellent agreement with our predictions in both peak position and spectral width, especially for configurations with smaller gyration radius. This observation supports previous interpretations from experimental55 and theoretical studies75, 76, suggesting that emission occurs from a broad ensemble of transient excited state geometries rather than from a single well-defined minimum. It is important to highlight that our theoretical spectrum reflects the recombination of the electron, since we describe the final electronic ground state as a closed shell (see Computational Methods). In contrast, the experimental measurements capture emission from the excited state of the hydrated electron without identifying the final state55, which could involve either recombination or hydrated electron in the ground state. Therefore, while the agreement between our results and the experimental data is encouraging, it should be considered preliminary. A more detailed investigation of these optical properties is necessary and lies beyond the scope of this study. However, our findings provide a valuable framework that could guide future experimental and theoretical efforts aimed at confirming and better understanding the emission behavior of the hydrated electron.
@@
Figure 6: Photon emission properties of the hydrated electron. Panel A: Histograms of the energy gap using all the frames of all the trajectories that undergo the PCET mechanism. Light blue bar is for all the frames with a gyration radius lower than 4 Å. Light purple bar is for all the frames with a gyration radius higher than 4 Å. Green and red solid lines are the experimental and fitted fluorescence spectra of the hydrated electron. Experimental data were digitalized from Ref. 55. Panel B: Time dependent evolution of the energy gap for a characteristic trajectory using ROKS method (green) and TDDFT (purple). The insets show the spin density and the particle orbital of the hydrated electron at 250 fs (denoted as a red point) for both ROKS and TDDFT methods, respectively. Panel C: Fluorescence spectra using the oscillator strength and energy gap for the last 300 fs for this trajectory using TDDFT.
@@ 3 Conclusions
We find that most excitations to the first band are localized on a single water molecule, but a significant fraction are delocalized over up to five molecules. This behavior, intrinsically linked to the asymmetry of the hydrogen-bond network, may explain the distinct photochemical outcomes observed at different energies within the first absorption band65, 66, 17. Our excited state molecular dynamics simulations reveal two distinct photophysical decay pathways: Hydrogen Atom Transfer (HAT) and Proton Coupled Electron Transfer (PCET). For the HAT mechanism, we predicted lifetimes and quantum yields are in excellent agreement with experimental data. Moreover, we observed the formation of the hydronium radical, a transient species previously proposed in theoretical studies but, to the best of our knowledge, not yet observed experimentally and reported here for the first time starting from photoexcited pure liquid water. This species has been suggested as a possible precursor to the hydrated electron71, 17. While our simulations, performed in the electronic excited state, can not exclude the possibility that it could release an electron in the ground electronic state, further studies are needed to confirm this pathway.
@@
In the PCET pathway, the hydrated electron emerges in the excited state. Our results further reveal that this localization is driven by ultrafast (<< 1 ps) collective rotational and translational motions of water molecules, particularly those participating in hydrogen bonding with the electron. In its early stages, the hydrated electron is highly delocalized. However, as localization proceeds, the gyration radius approaches values consistent with experimental measurements for the equilibrated hydrated electron in the electronic ground state. These findings indicate that a significant degree of electron localization together with collective waters rearrangements, occurs in the excited state before non-radiative decay. Moreover, we also observe the formation of hydroxyl radical–hydronium ion pairs, which was recently observed experimentally34. We investigated how their spatial separation influences both the gyration radius of the hydrated electron and the evolving energy gap during the excited-state dynamics. This correlation is particularly relevant in the early stages of the process, when diffusion is limited and the interactions between species are strongest.
@@
Finally, our work provides new insights into the broad fluorescence emission associated with the hydrated electron, as reported in experiments. We demonstrate a clear correlation between the emission energy and the degree of localization of the electron. This relationship offers a fresh perspective on the optical behavior of the hydrated electron. In recent studies77, 78, 79, fluorescence emission has been observed in different water solutions. Although different mechanisms have been proposed, the hydrated electron has not been considered as a contributing factor underlying these experimental observations.
@@
In conclusion, our theoretical studies here have provided fresh insights into the underlying mechanisms associated with on the most fundamental processes in the chemical physics and physical chemistry of water, namely how light interacts with liquid water and the subsequent photochemistry. Through a synergy of data-driven approaches, our approach lays the ground-work to study the birth of the excited-state hydrated electron in different bulk aqueous salt solutions77, 78, 79, at aqueous interfaces14, 80 and various biological contexts such as radiation-induced DNA damage4, 5. On the more methodological side, it is interesting to explore whether our simulations can be used to develop machine-learning approaches 81, 82, 83, 84, 85 to examine the excited-state dynamics of aqueous systems which would allow for longer simulations and more statistics. This is an area we are actively pursuing and will be the subject of a future study.
@@ 4.1 Ground-State Molecular Dynamics
We began by constructing a periodic simulation box containing 64 water molecules with a side length of 12.42 Å, corresponding to a density of 1 $g/cm^{3}$ at room temperature. An initial equilibration was performed using an empirical water model, SPC/E86 with the GROMACS software87 within the NVT ensemble for 100 ns. The canonical-sampling velocity-rescaling88 thermostat was applied with a coupling time of 0.1 ps.
From this classical equilibrated trajectory, we selected 100 configurations, separated by 1 ns, as initial conditions to perform ab initio quality ground-state molecular dynamics simulations with a machine-learning interatomic potential (MLIP). Specifically, we chose an MLIP recently trained and developed using the DeePMD-kit architecture89 on data computed using Density Functional Theory (DFT) with the SCAN0 functional90, which incorporates $10\%$ exact exchange to mitigate the self-interaction error. Prior studies90, 91, 92 have shown that SCAN0 provides an excellent description of the structural, dynamical, and electronic properties (such as IR spectra, electronic band structure, and band gap) of bulk liquid water compared to the original SCAN meta-GGA functional. Each of the 100 simulations conducted with the MLIP potential was run for 1 ns to generate well-converged and decorrelated initial conditions (positions and velocities), leveraging DeePMD-kit89 in combination with the LAMMPS package93.

### v2

@@ 2.3 H Atom Transfer Mechanism
Another frequently observed pathway is the formation of a species in which the dissociating hydrogen atom binds to a water molecule, forming the radical hydronium ($\mathrm{H_{3}O}^{\bullet}$), characterized by $CN_{H}\sim$1 and $CN_{O}\sim$3 (Figure 3C). We confirmed that this species is indeed $\mathrm{H_{3}O}^{\bullet}$ and not the conventional $\mathrm{H_{3}O}^{+}$ ion, since the electron remains associated with it. While the species $\mathrm{H_{3}O}^{\bullet}$ has been predicted by various theoretical models28, 56, 57, 70, our work is the first to demonstrate its formation following excitation by simulating the entire process from photoabsorption to relaxation in liquid water. According to our simulations, the radical $\mathrm{H_{3}O}^{\bullet}$ appears only transiently. It eventually undergoes further dissociation, where the hydrogen atom carrying the electron migrates into a region surrounded by water molecules. In these configurations, an OH bond from a nearby water molecule points toward the hydrogen atom (see Figure 3D) creating an unstable geometry ultimately leading to non-radiative decay to the ground-state. To the best of our knowledge the radical hydronium ($\mathrm{H_{3}O}^{\bullet}$) has not been observed experimentally through photoionization/photoabsorption of water perhaps in part due to the fact that it is formed fleetingly as a rare transition state-like motif.
@@ 2.4 Proton Coupled Electron Transfer Mechanism
Experimental studies have reported initial gyration radius of around 20-40 Å34, 67 immediately after photoabsorption. Although our method captures the same qualitative behavior, direct comparison with experiment is limited by the size of the simulation box (12.42 Å), which restricts the maximum observable extent of delocalization71. The final gyration radius we observe in the excited state is 2.7 Å, slightly larger than the experimentally reported value of 2.4 Å for the hydrated electron in the ground state72. Note that the decrease of the gyration radius from 6 to 2.4 Å incurs a decrease in the energy gap. Once the electron has localized, the system undergoes a non-radiative transition to the ground state, as indicated by the red points in Figure 4A.
As mentioned in the introduction, generating a free electron in bulk water using photon energies within the first absorption band requires some degree of nuclear rearrangement to release the electron from the molecule17. Earlier, we showed that this release occurs simultaneously with the dissociation of the OH bond in the excited water molecule, leading to the formation of a free electron, a hydroxyl radical ($\mathrm{HO}^{\bullet}$), and a proton ($\mathrm{H}^{+}$or $\mathrm{H_{3}O}^{+}$). We now move on to elucidating that the localization of the electron and the formation of the other species requires a collective process involving reorganization of the hydrogen bond network of the water molecules.
In Figure 4B we illustrate the coupling between the gyration radius against the angle formed between the vector connecting the oxygen and the center of the electron and the OH vector of the water molecule. For this analysis we only include the nearby water molecules involved in the first solvation shell of the electron, (see Analysis Protocols section for details). When the electron is delocalized, this angle spans a broad range from 0 to 120, reflecting the absence of a well-defined solvation shell. As the system relaxes, the angle narrows to below 30, consistent with the orientational preferences seen in hydrated electron solvation shells58, 71, 73. These results indicate that the structure of the hydrated electron forms within the excited-state itself, before the non-radiative decay consistent with recent experiments using sub-two-cycle visible-to-shortwave infrared pump–probe spectroscopy74. Analysis of the radial distribution function in the low gyration radius regime on the excited state (see Figure S9 in the Supplementary Information), where the hydrated electron is effectively localized, reveals that the first solvation shell comprises between 4-5 water molecules. This is consistent with previous findings for the hydrated electron in the ground electronic state45, 51, 25.
@@ 2.5 Coupled Translation and Rotational Solvent Dynamics Stabilizes Water-Mediated Photoproducts
In addition to these rotational dynamics, Figure 5B shows that localization also involves collective translational motion. Water molecules in the first solvation shell move on average by approximately 1.2 Å relative to their final positions, a displacement slightly larger than that of the more distant water molecules. Overall, our results show that both rotational and translational rearrangements of specific water molecules are necessary for the localization of the hydrated electron. These collective motions also cause a notable distortion of the hydrogen bond network, consistent with experimental findings from ionization studies in liquid water35.
As alluded to earlier, the creation of the excess electron after photoexcitation also leads to the formation of a hydronium ion ($\mathrm{H_{3}O}^{+}$) and a hydroxyl radical ($\mathrm{HO}^{\bullet}$). To investigate the behavior of these species during the excited-state dynamics, we present in Figure 5C the coupling between the gyration radius of the electron and the distance between the $\mathrm{HO}^{\bullet}$ and the $\mathrm{H_{3}O}^{+}$. Interestingly, our analysis reveals three distinct minima in the $\mathrm{HO}^{\bullet}$($\mathrm{H_{3}O}^{+}$) distance, particularly in the region of low gyration radius where the electron becomes localized. These minima occur at $\sim$2.6, 4.3, and 6.2 Å. Our simulations are in excellent agreement with very recent time-resolved electron diffraction experiments following the ionization of liquid water, where the corresponding minima were found at 2.4, 3.4, and 5.8 Å35. These three minima correspond to situations involving the initial formation of contact ion-radical pair and then the separation mediated by the solvent. Although previous theoretical studies have identified the $\mathrm{HO}^{\bullet}$($\mathrm{H_{3}O}^{+}$) pair by simulating ionized water with the electron removed75, 76, 77, our findings show for the first time, the creation of these correlated structures upon photoexcitation starting from the neutral bulk water.
The creation of ion-radical pairs at different distances has important implications on the optical properties. To probe this, we show in Figure 5D the energy gap as a function of the $\mathrm{HO}^{\bullet}$($\mathrm{H_{3}O}^{+}$) distance. For the contact ion-radical pair at 2.6 Å the average energy gap is $\sim$3 eV on average. As one moves towards the creation of the solvent-separated ion-radical pairs, the energy gap drops to $\sim$2.0 eV. This trend is also consistent with experimental observations, which demonstrated that the formation and spatial separation of the $\mathrm{HO}^{\bullet}$($\mathrm{H_{3}O}^{+}$) pair occur before the internal conversion associated with the $p\to s$ transition of the hydrated electron35.
@@ 2.6 Spontaneous Emission from Hydrated Excess Electrons
An important and still not fully understood property of the hydrated electron in liquid water is its photon emission following excitation. Understanding this emission can provide valuable insight into the nature of the excited state and the relaxation dynamics of the system. Although experimental studies are relatively scarce, Tauber and Mathies have explored this optical behavior through fluorescence and Raman spectroscopies58. Motivated by their findings, we now focus on analyzing the emission spectrum obtained from our simulations and examine how it depends on the gyration radius of the hydrated electron.
Figure 6A shows the experimental emission spectrum along with our calculated $S_{1}\to S_{0}$ energy gaps, evaluated for configurations with both large and small electron gyration radius. As the excess electron becomes more localized, transitioning from a larger to a smaller gyration radius, the energy gap decreases from an average of 2.4 to 1.3 eV (516 to 954 nm). Interestingly, the experimental emission spectrum is in excellent agreement with our predictions in both peak position and spectral width, especially for configurations with smaller gyration radius. This observation supports previous interpretations from experimental58 and theoretical studies78, 79, suggesting that emission occurs from a broad ensemble of transient excited state geometries rather than from a single well-defined minimum. It is important to highlight that our theoretical spectrum reflects the recombination of the electron, since we describe the final electronic ground state as a closed shell (see Computational Methods). In contrast, the experimental measurements capture emission from the excited state of the hydrated electron without identifying the final state58, which could involve either recombination or hydrated electron in the ground state. Therefore, while the agreement between our results and the experimental data is encouraging, it should be considered preliminary. A more detailed investigation of these optical properties is necessary and lies beyond the scope of this study. However, our findings provide a valuable framework that could guide future experimental and theoretical efforts aimed at confirming and better understanding the emission behavior of the hydrated electron.
@@
Figure 6: Photon emission properties of the hydrated electron. Panel A: Histograms of the energy gap using all the frames of all the trajectories that undergo the PCET mechanism. Light blue bar is for all the frames with a gyration radius lower than 4 Å. Light purple bar is for all the frames with a gyration radius higher than 4 Å. Green and red solid lines are the experimental and fitted fluorescence spectra of the hydrated electron. Experimental data were digitalized from Ref. 58. Panel B: Time dependent evolution of the energy gap for a characteristic trajectory using ROKS method (green) and TDDFT (purple). The insets show the spin density and the particle orbital of the hydrated electron at 250 fs (denoted as a red point) for both ROKS and TDDFT methods, respectively. Panel C: Fluorescence spectra using the oscillator strength and energy gap for the last 300 fs for this trajectory using TDDFT.
@@ 3 Conclusions
We find that most excitations to the first band are localized on a single water molecule, but a significant fraction are delocalized over up to five molecules. This behavior, intrinsically linked to the asymmetry of the hydrogen-bond network, may explain the distinct photochemical outcomes observed at different energies within the first absorption band68, 69, 18. Our excited state molecular dynamics simulations reveal two distinct photophysical decay pathways: Hydrogen Atom Transfer (HAT) and Proton Coupled Electron Transfer (PCET). For the HAT mechanism, we predicted lifetimes and quantum yields are in excellent agreement with experimental data. Moreover, we observed the formation of the hydronium radical, a transient species previously proposed in theoretical studies but, to the best of our knowledge, not yet observed experimentally and reported here for the first time starting from photoexcited pure liquid water. This species has been suggested as a possible precursor to the hydrated electron74, 18. While our simulations, performed in the electronic excited state, can not exclude the possibility that it could release an electron in the ground electronic state, further studies are needed to confirm this pathway.
@@
In the PCET pathway, the hydrated electron emerges in the excited state. Our results further reveal that this localization is driven by ultrafast (<< 1 ps) collective rotational and translational motions of water molecules, particularly those participating in hydrogen bonding with the electron. In its early stages, the hydrated electron is highly delocalized. However, as localization proceeds, the gyration radius approaches values consistent with experimental measurements for the equilibrated hydrated electron in the electronic ground state. These findings indicate that a significant degree of electron localization together with collective waters rearrangements, occurs in the excited state before non-radiative decay. Moreover, we also observe the formation of hydroxyl radical–hydronium ion pairs, which was recently observed experimentally35. We investigated how their spatial separation influences both the gyration radius of the hydrated electron and the evolving energy gap during the excited-state dynamics. This correlation is particularly relevant in the early stages of the process, when diffusion is limited and the interactions between species are strongest.
@@
Finally, our work provides new insights into the broad fluorescence emission associated with the hydrated electron, as reported in experiments. We demonstrate a clear correlation between the emission energy and the degree of localization of the electron. This relationship offers a fresh perspective on the optical behavior of the hydrated electron. In recent studies80, 81, 82, fluorescence emission has been observed in different water solutions. Although different mechanisms have been proposed, the hydrated electron has not been considered as a contributing factor underlying these experimental observations.
@@
In conclusion, our theoretical studies here have provided fresh insights into the underlying mechanisms associated with one the most fundamental processes in the chemical physics and physical chemistry of water, namely how light interacts with liquid water and the subsequent photochemistry. Through a synergy of data-driven approaches, our approach lays the ground-work to study the birth of the excited-state hydrated electron in different bulk aqueous salt solutions80, 81, 82, at aqueous interfaces14, 83 and various biological contexts such as radiation-induced DNA damage4, 5. On the more methodological side, it is interesting to explore whether our simulations can be used to develop machine-learning approaches 84, 85, 86, 87, 88 to examine the excited-state dynamics of aqueous systems which would allow for expanding the sampling and statistics. This is an area we are actively pursuing and will be the subject of a future study.
@@ 4.1 Ground-State Molecular Dynamics
We began by constructing a periodic simulation box containing 64 water molecules with a side length of 12.42 Å, corresponding to a density of 1 $g/cm^{3}$ at room temperature. An initial equilibration was performed using an empirical water model, SPC/E89 with the GROMACS software90 within the NVT ensemble for 100 ns. The canonical-sampling velocity-rescaling91 thermostat was applied with a coupling time of 0.1 ps.
From this classical equilibrated trajectory, we selected 100 configurations, separated by 1 ns, as initial conditions to perform ab initio quality ground-state molecular dynamics simulations with a machine-learning interatomic potential (MLIP). Specifically, we chose an MLIP recently trained and developed using the DeePMD-kit architecture92 on data computed using Density Functional Theory (DFT) with the SCAN0 functional93, which incorporates $10\%$ exact exchange to mitigate the self-interaction error. Prior studies93 have shown that SCAN0 provides an excellent description of the structural, dynamical, and electronic properties (such as IR spectra, electronic band structure, and band gap) of bulk liquid water compared to the original SCAN meta-GGA functional94, 95. Each of the 100 simulations conducted with the MLIP potential was run for 1 ns to generate well-converged and decorrelated initial conditions (positions and velocities), leveraging DeePMD-kit92 in combination with the LAMMPS package96.